What is the budgeted gross profit for June?


Thomas Corporation recorded sales of $119,163 during March. Management expects sales to increase 5% in April, another 2% in May, and another 10% in June. cost of goods sold is expected to be 61% of sales. What is the budgeted gross profit for June?
Answer #1

Sales in June= Sales in March*(1+5%)*(1+2%)*(1+10%)=119163*1.05*1.02*1.1=$140385.93


COGS=Sales*68%=$85635.41


Gross Prfoit=Sales-COGS=$54750.52


answered by: Brennands